OpenStudy (brydenwright):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (mathmale):

Bryden, could you find and post the formula that involves "cos theta" and the components of vectors u and v? You are given all of these comps, and thus have enuf info by which to find cos theta, the angle between the given vectors. Final step is to use the inverse cosine function (on your calculator, perhaps?) to determine the angle, theta. Questions?

OpenStudy (mathmate):

You can use the same formula as in your previous post: http://openstudy.com/study#/updates/57c8dee5e4b0948881943490 If you could use an example, here it is: Find the angle between the given vectors to the nearest tenth of a degree. u = <-3, 5>, v = <7, -2> u.v=(-3)(7)+(5)(-2)=-21-10=-31 ||u||=sqrt((-3)^2+5^2)=sqrt(9+25)=sqrt(34) ||v||=sqrt(7^2+(-2)^2)=sqrt(49+4)=sqrt(53) \(cos(\theta)=-31/(\sqrt{34}\sqrt{53})\)=-0.73027 so \(\theta\)=\(cos^{-1}(-.73027)=136.91 or 223.09\) degrees. The previous link also contains a link to a detailed example: http://www.wikihow.com/Find-the-Angle-Between-Two-Vectors